# Electric field induction in quark-gluon plasma due to thermoelectric effects

### v1

@@ Abstract
Relativistic heavy-ion collisions produce quark-gluon plasma (QGP), which is locally thermalized. QGP with higher thermal conductivity advances towards global thermalization. Being made of electrically charged particles (quarks), QGP exhibits interesting thermoelectric phenomena during its evolution, resulting in the induction of an electric field in the medium. For the first time, we estimate the induced electric field in the QGP due to thermoelectric effect. This can be seen even in the QGP produced in the head-on collisions. The Seebeck coefficient is essential in determining the induced field. However, spectator current can produce a magnetic field in peripheral heavy-ion collisions. This breaks the isotropy of the thermoelectric coefficient matrix and introduces magneto-Seebeck and Nernst coefficients that contribute to the induced electric field in peripheral collisions. We have taken care of the temperature evolution of QGP with different hydrodynamic cooling rates to calculate the transport coefficients. The induced electric field is estimated with the cooling rate obtained from Gubser hydrodynamic flow. We estimated the space-time profile of the induced field and found that it is zero at the center and increases as we go away from the center. At the early time of evolution, the electric field can reach a maximum value of $eE\approx 1~{}m_{\pi}^{2}$, and the strength decreases in time. Moreover, we estimated the transport coefficients in the presence of the external time-varying magnetic field. The effect of the intensity and decay parameter of the magnetic field on the induced electric field is also explored.
arxiv: $0$
1 Introduction
The success of heavy-ion collision experiments is not limited to the discovery of phenomena of strong interaction at very high energy but also contributes to understanding fundamental physics applicable to many different areas, such as condensed matter physics. The quark-gluon plasma (QGP), which is a deconfined and locally thermalized state of quarks and gluons, produced in relativistic heavy-ion collisions (RHICs), undergoes many phases during its evolution BREWER2021136189 ; Busza:2018rrf . Some phenomena occurred during the evolution Elfner:2022iae ; PhysRevC.106.014905 significantly impact final state observables to be distinguished, and some might not yet be discovered LIAO201699 . It is found that relativistic dissipative hydrodynamics is the most effective theory that can successfully explain final state observables such as spectra and flow PhysRevLett.87.182301 ; PhysRevC.77.054901 ; article . The thermodynamic PhysRevD.100.076009 ; PhysRevC.99.035210 ; Goswami:2023eol and transport Gavin:1985ph ; PhysRevD.100.114004 ; PhysRevD.102.016016 ; PhysRevD.106.034008 ; PhysRevD.108.094007 properties of nearly baryon-free QGP matter created in the Large Hadron Collider (LHC) and that of QGP with finite baryon-chemical potential created in Relativistic heavy-ion Collider (RHIC) are explored very well. Moreover, going from head-on to peripheral collisions, an intense transient electromagnetic field is created that can be understood from laws of classical electrodynamics Tuchin:2015oka ; PhysRevD.90.034018 ; PhysRevLett.125.022301 ; KHARZEEV2008227 ; PhysRevX.14.011028 . The change in energy and centrality not only modifies the thermodynamic equation of state (EoS) Kumar2024 ; Brandt2023 ; Borsanyi2012 ; Bali2014 of the created medium but many different phenomena, such as chiral magnetic effect KHARZEEV2008227 ; PhysRevD.97.056024 , magnetic and inverse magnetic catalysis PhysRevLett.73.3499 ; PhysRevD.55.6504 , and plausible transition through the quantum chromodynamics (QCD) critical point PhysRevD.85.091901 ; Endrodi2011 .
@@ 1 Introduction
The paper is organized in the following manner. First, we briefly derive the cooling rates for Bjorken and Gubser flow in section (2). In section (3), we derive the Seebeck coefficient for the head-on collision or $eB=0$ case. Then, we estimate the Seebeck coefficient and induced electric field with QGP EoS. Then, we extend our study to peripheral collision in section (3) where the initial magnetic field is non-zero. We calculate the magneto-Seebeck and Nernst coefficients for the anisotropic system and calculate the induced electric field for the same in section (4). Finally, in section (5), we have summarized our study with possible outlook. Detailed calculations for thermoelectric coefficients are given in the appendix.
2 Relativistic hydrodynamics and cooling rates
The relativistic hydrodynamics is a successful macroscopic theory to understand the dynamics of relativistic fluid created in RHICs, namely QGP Rischke:1995mt ; Okamoto2016 . Hydrodynamics and the correct EoS of a medium will provide us with a detailed dynamical evolution of the medium and its cooling rate or space-time-dependent temperature profile. Here, to estimate thermoelectric coefficients and induced electric field, an analytical expression of cooling rate is required, for which we use Bjorken and Gubser flow PhysRevD.27.140 ; PhysRevD.82.085027 .
The symmetric energy-momentum tensor ($T^{\mu\nu}$) of a fluid can be expressed in terms of hydrodynamic degrees of freedom as PhysRevC.103.014905 ,
@@ 2 Relativistic hydrodynamics and cooling rates
leads to the continuity equation PhysRevLett.99.172301 . One can obtain the required cooling rate by solving Eq. (2) with the relevant flow profile and EoS.
2.1 Bjorken flow
The Bjorken flow represents a boost invariant longitudinal expansion of the medium created in heavy-ion collisions PhysRevD.27.140 . Along with longitudinal boost invariant, Bjorken considered translation and rotational invariance symmetry in the transverse plane, which leads to the velocity profile
@@ 2.1.1 Ideal hydrodynamics at $B=0$
Solving the hydrodynamic Eq. (2) by using Bjorken flow Eq. (3) and ideal energy-momentum tensor Eq. (4) we get PhysRevD.27.140
@@
2.1.2 Ideal magnetohydrodynamics
In peripheral RHICs, a huge magnetic field is created, which can magnetize the QGP. The energy-momentum tensor for an ideal magnetized fluid can be expressed as ROY201545 ; PhysRevD.109.014018 ; Panda2021
@@ 2.1.2 Ideal magnetohydrodynamics
Depending on the medium’s electrical conductivity, the magnetic field created during the collision will decay with time. Here, we consider an exponential decay profile Satow:2014lia
@@
2.2 Gubser flow
Gubser generalizes the Bjorken flow to include radial expansion of the medium and longitudinal boost-invariant expansion PhysRevD.82.085027 . Working with rapidity or the Milne coordinate system is convenient in heavy-ion collisions. Milne coordinates ($\tau,\eta,r,\phi$), can be related to the Cartesian coordinates $(t,x,y,z)$ as,
$\tau=\sqrt{t^{2}-z^{2}},~{}~{}~{}~{}\eta=arctanh\frac{z}{t},~{}~{}~{}~{}~{}r=\sqrt{x^{2}+y^{2}},~{}~{}~{}~{}\phi=arctan\frac{y}{x}.$ (11)
@@ 2.2 Gubser flow
$u^{\tau}=\cosh{\kappa}=\gamma_{r},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\frac{u^{r}}{u^{\tau}}=\tanh{\kappa}=v_{r},~{}~{}~{}~{}~{}~{}u^{\eta}=u^{\phi}=0.$ (12)
@@
Where Lorentz factor $\gamma_{r}$ = $\frac{1}{\sqrt{1-v_{r}^{2}}}$ and $v_{r}$ = $\sqrt{\vec{v}_{x}^{2}+\vec{v}_{y}^{2}}$, $v_{r}$ is known as transverse velocity. $\kappa$ is parameterized as $\kappa(\tau,r)={arctanh}(\frac{2q^{2}\tau r}{1+q^{2}\tau^{2}+q^{2}r^{2}})$. Here, we take the case of the viscous medium while preserving the conformal invariance of the theory, that is, $p=\epsilon/3$ and shear viscosity $\eta$= $H_{0}\epsilon^{3/4}$. $H_{0}$ is a dimensionless quantity. Using the Gubser flow, we can solve the hydrodynamic equation for viscous medium to obtain the temperature profile as PhysRevD.82.085027 ,
$T$ $={1\over\tau f_{*}^{1/4}}\left[{\frac{{T}_{0}}{(1+g^{2})^{1/3}}}+\frac{{\rm H}_{0}g}{\sqrt{1+g^{2}}}\left\{1-(1+g^{2})^{1/6}{}_{2}F_{1}\left({1\over 2},{1\over 6};{3\over 2};-g^{2}\right)\right\}\right]~{}.$ (13)
@@
Where ${T}_{0}$ is an integration constant and ${}_{2}F_{1}$ denotes a hypergeometric function with $g={1-q^{2}\tau^{2}+q^{2}r^{2}\over 2q\tau}$. For numerical estimations, we used semi-realistic numbers for a central gold-gold collision at $\sqrt{s_{\rm NN}}=200\,{\rm GeV}$ are $\hat{T}_{0}=5.55$ and ${\rm H}_{0}=0.33$ if we choose $4.31/q=$ 4.3 fm.
@@
Figure 1: Temperature ($T$) as a function of proper time ($\tau$) for ideal hydrodynamics, ideal magnetohydrodynamics (MHD) with two decay parameters at $\tau_{B}=3,~{}7$ fm, and Gubser flow at $r=0,~{}3$ fm.
@@
In Fig. (1), we have plotted the cooling rate (temperature as a function of proper time) for all the three hydrodynamic cases discussed above. Note that the initial conditions are free to adjust within the Heisenberg uncertainty principle. The cooling rates are represented as follows. The ideal hydrodynamics at $B=0$ with a black solid line, ideal magnetohydrodynamics (MHD) at $\tau_{B}=3,~{}7$ fm with the red dashed-triple dot line and green dash-dot line, respectively, and for Gubser flow at $r=0,~{}3$ fm by magenta dashed and blue dashed line, respectively. Interestingly, the temperature in Gubser flow falls faster due to the presence of transverse flow even at the finite viscosity of the medium. The cooling rates plotted here will be helpful in the analysis of results obtained in the subsequent sections.
3 Head-on collision
@@
In this section, we calculate the thermoelectric coefficient for the QGP medium created in a head-on collision. Due to the absence of spectator current, no magnetic field is expected to be created in the head-on collisions. Here, we will show that an electric field can be created or induced in the medium due to the thermoelectric effect, even though there is no initial magnetic field during the collision.
3.1 Seebeck coefficient in evolving QGP
@@
QGP created in RHICs cools down rapidly as it goes through space-time evolution with rapid expansion. This rapid cooling affects the medium’s transport coefficients, such as thermal conductivity and thermoelectric coefficients. In all the earlier studies of thermoelectric coefficients Das2021 ; Abhishek:2020wjm ; Kurian:2021zyb ; PhysRevD.102.096011 , a static picture of QGP is considered where the effect of temperature evolution was not taken into account, which is a more realistic scenario. In Ref. PhysRevD.109.014018 , for the first time, we have estimated the effect of temperature evolution on the thermal conductivity of QGP. Here, we will calculate the thermoelectric coefficient for evolving QGP using a similar approach of Ref. PhysRevD.109.014018 . We will also show the difference between the expressions obtained considering a static QGP.
@@
where $\pm$ stands for fermion and boson, respectively. The total single-particle distribution function ($f_{i}$) for a system slightly out of equilibrium ($\delta f_{i}$) can be written as $f_{i}=f^{0}_{i}+\delta f_{i}$. When a system is out of equilibrium, net non-zero currents arise in the system. In kinetic theory, electric current density for such a system can be expressed as PhysRevD.108.094007
@@
Here, $q_{i}$ is the electric charge, and $g_{i}$ is the degeneracy of the ith species particles. Similarly, the microscopic definition of heat flow PhysRevD.108.094007 ; PhysRevD.109.014018 ,
@@
Where $h$ is the total enthalpy per particle. To find the expression of $\delta f_{i}$, we solve the Boltzmann transport equation (BTE) with the help of relaxation time approximation (RTA) PhysRevD.102.016016
@@
where $\tau^{i}_{R}$ is the relaxation time of the particle. $\vec{E}$ represents the external electric field. Here, the deviation of the distribution function from equilibrium is driven by the electric field and temperature gradients. With leading order contribution, we can consider an ansatz of $\delta f_{i}$ as PhysRevD.104.094037
@@
The unknown coefficients $\alpha_{j}$ ($j=1,2,3$) determine the strength of the respective gradient force fields driving the system away from equilibrium. The coefficient $\alpha_{3}$ corresponds to the term that arises due to the evolving picture of the medium. For the case of a static picture, where the effects of cooling are not considered, only $\alpha_{1}$ and $\alpha_{2}$ contribute, whereas $\alpha_{3}$ vanishes in Eq. (19) as in Ref. Das2021 ; Kurian:2021zyb . We can solve Eq. (17) using Eqs. (18) and (19), and get the expression of $\alpha_{j}$’s. After getting the expressions for $\alpha_{j}$’s, the simplified form of $\delta f_{i}$ is (see Appendix A),
$\delta f_{i}$ $=-q_{i}\tau_{R}^{i}(\vec{v_{i}}.\vec{E})+\frac{(\omega_{i}-{\rm b}_{i}h)\tau_{R}^{i}}{T}\Big{[}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{T}\Big{)}-\tau_{R}^{i}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{\dot{T}}\Big{)}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (20)
@@
$j^{l}$ $=\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}|\vec{k}_{i}|}{(2\pi)^{3}}v_{i}^{2}\tau_{R}^{i}\Big{[}-q_{i}E^{l}+\frac{(\omega_{i}-{\rm b}_{i}h)}{T}\Big{\{}\frac{\partial{T}}{\partial x^{l}}-\tau_{R}^{i}\frac{\partial{\dot{T}}}{\partial x^{l}}\Big{\}}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (21)
@@
We introduce the following integrals to write Eqs. (3.1) in compact form as
@@
$\bm{\sigma}=\begin{pmatrix}\sigma_{e}&0&0\\ 0&\sigma_{e}&0\\ 0&0&\sigma_{e}\\ \end{pmatrix},~{}\hskip 28.45274pt\mathbb{E}=\begin{pmatrix}E_{x}\\ E_{y}\\ E_{z}\end{pmatrix},~{}$
$\mathbb{L}=\frac{1}{T}\begin{pmatrix}L_{23}&0&0\\ 0&L_{23}&0\\ 0&0&L_{23}\\ \end{pmatrix},~{}\mathbb{X}=\begin{pmatrix}\frac{dT}{dx}\vspace{2mm}\\ \frac{dT}{dy}\vspace{2mm}\\ \frac{dT}{dz}\\ \end{pmatrix}.$
@@
3.1.1 Analytical solution for massless non-interacting system
Here, we address the analytic solution of Seebeck coefficient Eq. (27) for massless ($\omega_{i}=k_{i}$) and non-interacting system. In this case, the Seebeck coefficient can be solved analytically using the Riemann zeta function $\zeta(s)$ and Dirichlet eta $\eta(s)$ function:

### v2

@@ Abstract
Relativistic heavy-ion collisions produce quark-gluon plasma (QGP), which is locally thermalized. Due to electrically charged particles (quarks), QGP exhibits interesting thermoelectric phenomena during its evolution, resulting in an electromagnetic (EM) field in the medium. In this study, for the first time, we estimate the induced electric field in QGP due to the thermoelectric effect. This phenomenon can induce an EM field even in QGP produced by the head-on heavy-ion collision. In peripheral heavy-ion collisions, the presence of a spectator current generates a transient magnetic field at the early stage, which disrupts the isotropy of the induced electric field. For the numerical estimation, we use a quasiparticle-based model that incorporates the lattice quantum chromodynamics equation of state for QGP. The induced electric field is estimated with cooling rates derived from Gubser hydrodynamic flow. Thermoelectric coefficients such as Seebeck, magneto-Seebeck, and Nernst coefficients play a crucial role in determining the induced field. Additionally, we account for the temperature evolution of QGP using different hydrodynamic cooling rates to calculate the transport coefficients. We also estimate the transport coefficients and the induced electric field in the presence of an external time-varying magnetic field, including the quantum effect of Landau quantization, and explore the effects of the intensity and decay parameter of the magnetic field on the induced electric field. Our findings reveal that the space-time profile of the induced electric field is zero at the center and increases as we go away from the center. During the early stages of evolution, the electric field can reach a maximum value of $eE\approx 1~{}m_{\pi}^{2}$, decreasing in strength over time.
I Introduction
@@
In the present work, we estimate the induced electric field in QGP using the thermoelectric effect. The creation of electric potential due to temperature gradient or vice versa is known as thermoelectric effect [37]. In baryonic QGP, the heat flow is governed by the baryon current carried by the constituent quarks. As the quarks are electrically charged, a net nonzero electric potential develops, resulting in an electric field. This field is estimated using thermoelectric transport coefficients, such as Seebeck, magneto-Seebeck, and Nernst coefficients. In RHICs, these coefficients are studies for QGP [38, 39] and hadronic matter [40, 41] in zero and finite external magnetic field cases. However, none of these investigations estimated the electromagnetic field created in QGP due to thermoelectric processes. In this work, we assessed the induced electric field and looked at the dynamic image of QGP, considering the influence of temperature evolution (or cooling). In a recent study [42], we demonstrated the significance of temperature evolution in QGP observables like elliptic flow. It is essential to mention that unlike the EM field produced in peripheral heavy-ion collisions due to spectators, this field can be induced in QGP created in head-on collisions where a nonzero baryon chemical potential is expected. We also estimate the thermoelectric effect-induced electric field in a peripheral collision where a strong transient-external magnetic field in the early stage is expected to affect the QGP significantly. Where the quantum modification of energy states due to Landau quantization is incorporated. This is the first time an electric field generated solely by the thermoelectric effect has been estimated for QGP. The created spatiotemporal electric field can generate a spatiotemporal magnetic field, which is also discussed, along with possible limitations.
The paper is organized in the following manner. First, we briefly derive the cooling rates for Bjorken and Gubser flow in Sec.II. In Sec.III, we derive the Seebeck coefficient for the head-on collision or $eB=0$ case. Then, we estimate the Seebeck coefficient and induced electric field with QGP EoS. Then, we extend our study to peripheral collision in Sec.III where the initial magnetic field is nonzero. We calculate the magneto-Seebeck and Nernst coefficients for the anisotropic system and calculate the induced electric field for the same in Sec.IV. Finally, in Sec.V, we have summarized our study with a possible outlook. Detailed calculations for thermoelectric coefficients are given in the appendix.
II Relativistic hydrodynamics and cooling rates
The relativistic hydrodynamics is a successful macroscopic theory to understand the dynamics of relativistic fluid created in RHICs, namely QGP [43, 44]. Hydrodynamics and the correct EoS of a medium will provide us with a detailed dynamical evolution of the medium and its cooling rate or space-time-dependent temperature profile. Here, to estimate thermoelectric coefficients and induced electric field, an analytical expression of cooling rate is required, for which we use Bjorken and Gubser flow [45, 46].
The symmetric energy-momentum tensor ($T^{\mu\nu}$) of a fluid can be expressed in terms of hydrodynamic degrees of freedom as [47],
@@ II Relativistic hydrodynamics and cooling rates
leads to the continuity equation [48]. One can obtain the required cooling rate by solving Eq. (2) with the relevant flow profile and EoS.
II.1 Bjorken flow
The Bjorken flow represents a boost invariant longitudinal expansion of the medium created in heavy-ion collisions [45]. Along with longitudinal boost invariant, Bjorken considered translation and rotational invariance symmetry in the transverse plane, which leads to the velocity profile
@@ II.1.1 Ideal hydrodynamics at $B=0$
Solving the hydrodynamic Eq. (2) by using Bjorken flow Eq. (3) and ideal energy-momentum tensor Eq. (4) we get [45]
@@
II.1.2 Ideal magnetohydrodynamics
In peripheral RHICs, a huge magnetic field is created, which can magnetize the QGP. The energy-momentum tensor for an ideal magnetized fluid can be expressed as [49, 42, 50]
@@ II.1.2 Ideal magnetohydrodynamics
Depending on the medium’s electrical conductivity, the magnetic field created during the collision will decay with time. Here, we consider an exponential decay profile [51]
@@
II.2 Gubser flow
Gubser generalizes the Bjorken flow to include radial expansion of the medium and longitudinal boost-invariant expansion [46]. Working with rapidity or the Milne coordinate system is convenient in heavy-ion collisions. Milne coordinates ($\tau,\eta,r,\phi$), can be related to the Cartesian coordinates $(t,x,y,z)$ as,
$\tau=\sqrt{t^{2}-z^{2}},~{}~{}~{}~{}\eta=arctanh\frac{z}{t},$
$r=\sqrt{x^{2}+y^{2}},~{}~{}~{}~{}\phi=arctan\frac{y}{x}.$ (11)
@@ II.2 Gubser flow
$u^{\tau}=\cosh{\kappa}=\gamma_{r},~{}~{}~{}\frac{u^{r}}{u^{\tau}}=\tanh{\kappa}=v_{r},~{}~{}~{}u^{\eta}=u^{\phi}=0.$ (12)
@@
Where Lorentz factor $\gamma_{r}$ = $\frac{1}{\sqrt{1-v_{r}^{2}}}$ and $v_{r}$ = $\sqrt{\vec{v}_{x}^{2}+\vec{v}_{y}^{2}}$, $v_{r}$ is known as transverse velocity. $\kappa$ is parametrized as $\kappa(\tau,r)={arctanh}(\frac{2q^{2}\tau r}{1+q^{2}\tau^{2}+q^{2}r^{2}})$. Here, we take the case of the viscous medium while preserving the conformal invariance of the theory, that is, $p=\epsilon/3$ and shear viscosity $\eta$= $H_{0}\epsilon^{3/4}$. $H_{0}$ is a dimensionless quantity. Using the Gubser flow, we can solve the hydrodynamic equation for viscous medium to obtain the temperature profile as [46],
$T$ $={1\over\tau f_{*}^{1/4}}\Big{[}{\frac{{T}_{0}}{(1+g^{2})^{1/3}}}+\frac{{\rm H}_{0}g}{\sqrt{1+g^{2}}}\Big{\{}1-(1+g^{2})^{1/6}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{}_{2}F_{1}\left({1\over 2},{1\over 6};{3\over 2};-g^{2}\right)\Big{\}}\Big{]}~{}.$ (13)
@@
Where ${T}_{0}$ is an integration constant and ${}_{2}F_{1}$ denotes a hypergeometric function with $g={1-q^{2}\tau^{2}+q^{2}r^{2}\over 2q\tau}$. For numerical estimations, we used semirealistic numbers for a central gold-gold collision at $\sqrt{s_{\rm NN}}=200\,{\rm GeV}$ are $\hat{T}_{0}=5.55$ and ${\rm H}_{0}=0.33$ if we choose $4.31/q=$ 4.3 fm.
@@
Figure 1: Temperature ($T$) as a function of proper time ($\tau$) for ideal hydrodynamics, ideal magnetohydrodynamics (MHD) for $eB_{0}=5~{}m_{\pi}^{2}$ with two decay parameters at $\tau_{B}=3,~{}7$ fm, and Gubser flow at $r=0,~{}3$ fm.
@@
In Fig. (1), we have plotted the cooling rate (temperature as a function of proper time) for all the three hydrodynamic cases discussed above. Note that the initial conditions are free to adjust within the Heisenberg uncertainty principle. The cooling rates are represented as follows. The ideal hydrodynamics at $B=0$ with a black solid line, ideal magnetohydrodynamics (MHD) at $\tau_{B}=3,~{}7$ fm with the red dashed-triple dot line and green dash-dot line, respectively, and for Gubser flow at $r=0,~{}3$ fm by magenta dashed and blue dashed line, respectively. Interestingly, the temperature in Gubser flow falls faster due to the presence of transverse flow even at the finite viscosity of the medium. The cooling rates plotted here will be helpful in the analysis of results obtained in the subsequent sections. Note that the magnetic decay parameter values considered here are high and unrealistic in HIC. In principle, it depends on many parameters, such as QGP electrical conductivity and impact parameters, and can be determined through MHD simulation. However, considering such a high parameter value allows us to distinguish the decay effect in the plots, which gives us a qualitative dependency on the decay parameter. For more insight into the lifetime of the electromagnetic field in QGP due to spectator current, see Refs. [52, 53, 54, 51, 55].
III Electric field induction in head-on collisions
@@
This section calculates the thermoelectric coefficient and induced electric field for the QGP medium created in a head-on collision. Due to the absence of spectator current, no magnetic field is expected to be created in the head-on collisions. Here, we will show that an electric field can be induced in the medium due to the thermoelectric effect, even though there is no initial EM field during the collision.
III.1 Seebeck coefficient in evolving QGP
@@
QGP created in RHICs cools down rapidly as it goes through space-time evolution with rapid expansion. This rapid cooling affects the medium’s transport coefficients, such as thermal conductivity and thermoelectric coefficients. In all the earlier studies of thermoelectric coefficients [38, 56, 57, 39], a static picture of QGP is considered where the local temperature gradients are present i.e. $\frac{\partial{T}}{\partial x^{k}}\neq$ 0, but the effects of cooling are not considered during its evolution i.e. $\frac{\partial{T}}{\partial t}$ = 0, which is a more realistic scenario. In Ref. [42], for the first time, we have estimated the effect of temperature evolution on the thermal conductivity of QGP. Here, we will calculate the thermoelectric coefficient for evolving QGP using a similar approach of Ref. [42]. We will also show the difference between the expressions obtained considering a static QGP.
@@
where $\pm$ stands for fermion and boson, respectively. The total single-particle distribution function ($f_{i}$) for a system slightly out of equilibrium ($\delta f_{i}$) can be written as $f_{i}=f^{0}_{i}+\delta f_{i}$. When a system is out of equilibrium, net nonzero currents arise in the system. In kinetic theory, electric current density for such a system can be expressed as [17]
@@
Here, $q_{i}$ is the electric charge, and $g_{i}$ is the degeneracy of the ith species particles. Similarly, the microscopic definition of heat flow [17, 42],
@@
Where $h$ is the total enthalpy per particle. To find the expression of $\delta f_{i}$, we solve the Boltzmann transport equation (BTE) with the help of relaxation time approximation (RTA) [15]
@@
where $\tau^{i}_{R}$ is the relaxation time of the particle. $\vec{E}$ represents the external electric field. Here, the deviation of the distribution function from equilibrium is driven by the electric field and temperature gradients. With leading order contribution, we can consider an ansatz of $\delta f_{i}$ as [58]
@@
The unknown coefficients $\alpha_{j}$ ($j=1,2,3$) determine the strength of the respective gradient force fields driving the system away from equilibrium. The coefficient $\alpha_{3}$ corresponds to the term that arises due to the evolving picture of the medium. For the case of a static picture, where the effects of cooling are not considered, only $\alpha_{1}$ and $\alpha_{2}$ contribute, whereas $\alpha_{3}$ vanishes in Eq. (19) as in Ref. [38, 57]. We can solve Eq. (17) using Eqs. (18) and (19), and get the expression of $\alpha_{j}$’s. After getting the expressions for $\alpha_{j}$’s, the simplified form of $\delta f_{i}$ is (see Appendix A),
$\delta f_{i}$ $=-q_{i}\tau_{R}^{i}(\vec{v_{i}}.\vec{E})+\frac{(\omega_{i}-{\rm b}_{i}h)\tau_{R}^{i}}{T}\Big{[}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{T}\Big{)}$
$-\tau_{R}^{i}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{\dot{T}}\Big{)}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (20)
@@
$j^{l}=$ $\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}|\vec{k}_{i}|}{(2\pi)^{3}}v_{i}^{2}\tau_{R}^{i}\Big{[}-q_{i}E^{l}+\frac{(\omega_{i}-{\rm b}_{i}h)}{T}$
$\Big{\{}\frac{\partial{T}}{\partial x^{l}}-\tau_{R}^{i}\frac{\partial{\dot{T}}}{\partial x^{l}}\Big{\}}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (21)
@@
We introduce the following integrals to write Eqs. (III.1) in compact form as
@@
$\bm{\sigma}=\begin{pmatrix}\sigma_{e}&0&0\\ 0&\sigma_{e}&0\\ 0&0&\sigma_{e}\\ \end{pmatrix},~{}\hskip 28.45274pt\mathbf{E}=\begin{pmatrix}E_{x}\\ E_{y}\\ E_{z}\end{pmatrix},~{}$
$\mathbf{L}=\frac{1}{T}\begin{pmatrix}L_{23}&0&0\\ 0&L_{23}&0\\ 0&0&L_{23}\\ \end{pmatrix},~{}\mathbf{X}=\begin{pmatrix}\frac{dT}{dx}\vspace{2mm}\\ \frac{dT}{dy}\vspace{2mm}\\ \frac{dT}{dz}\\ \end{pmatrix}.$
@@
III.1.1 Analytical solution for massless noninteracting system
Here, we address the analytic solution of Seebeck coefficient Eq. (27) for massless ($\omega_{i}=k_{i}$) and noninteracting system. In this case, the Seebeck coefficient can be solved analytically using the Riemann zeta function $\zeta(s)$ and Dirichlet eta $\eta(s)$ function: